A coffee shop is evaluating the effect of cutting the price of premium coffee from R40 to R38 a cup. Current sales are 3 800 cups per year. The firm believes the price increase will result in selling 200 additional cups. It also expects increased sales of chocolate cake, as a result of the coffee promotion - from 200 to 220 cakes per year

1.what is the arc cross elasticity of demand for cakes with respect to coffee
2. is the firm's expectation of selling more chocolate cakes justified? explain
1
Expert's answer
2019-08-23T15:36:19-0400

1. The arc cross elasticity of demand for cakes with respect to coffee is:

Ed = (220 - 200)/(38 - 40)×(40 + 38)/(220 + 200) = -20/2×78/420 = -1.86.

2. The firm's expectation of selling more chocolate cakes is justified, because these goods are complements.